Use the fact that the earth is is 3960 miles and 1 mile is 5280 feet answer the following question.
A tower of a submarine is approximately 15 feet above sea level. How far can a person see from the top of the tower?

Answer by RAY100(1637)

Radius of the Earth is approx 3960 mi
.
If a tower is 15 ft above this, the distance is 15 ft /5280 ft/mi = .00284 mi,,,,,and the distance to the ctr of Earth is 3960 + .00284 mi
.
If we construct a right triangle, one vertex at the top of the tower,
one at the center of the Earth, and another at the horizon point from the tower,,,,we find a right triangle, one leg 3960 mi, the hypotenuse = 3960.00284 mi, and the other leg the distance to the horizon. ( this leg is tan to curvature of Earth,,hence 90 degree angle)
.
using the Pythagorean Theorem ,,c^2 = a^2 +b^2,
.
(3960,00284)^2 = d^2 + 39606^2
.
d= 4.7434 miles,,,,,,distance to horizon from tower.
.